40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 27 20:41:21 2017, rana, gautam, Update, PSL, PMC DAQ assay for feed-forward integration TEK00000.PNG20170327_194931.jpg20170327_204554.jpg
    Reply  Wed Mar 29 11:41:19 2017, gautam, Update, PSL, PMC DAQ assay for feed-forward integration C1ALS_PSLnamespace.pngPMCcheckout.pdfD980352-A-40m_151119.pdf
       Reply  Thu Mar 30 00:16:09 2017, gautam, Update, PSL, PSL NPRO PZT calibration 
          Reply  Mon Apr 3 17:25:13 2017, gautam, Update, PSL, PSL NPRO PZT calibration PDHerr.pdfNPROcalib.pdf
             Reply  Tue Apr 4 17:27:58 2017, rana, Update, PSL, PSL NPRO PZT calibration 
       Reply  Thu Mar 30 18:11:01 2017, rana, Update, PSL, PMC DAQ assay for feed-forward integration 
          Reply  Mon Apr 3 23:07:09 2017, gautam, Update, PSL, PMC DAQ assay for feed-forward integration 
             Reply  Tue Apr 11 00:38:37 2017, gautam, Update, PSL, PMC demod moved off servo board PMC_Ctrl_spec.pdf
                Reply  Wed Apr 12 00:36:53 2017, gautam, Update, PSL, PMC demod moved off servo board PMC_digitalSpec.pdfPMC_DAQ_spectra.pdf
                   Reply  Tue Apr 18 01:01:03 2017, gautam, Update, PSL, PMC OLTF measured, DAQ channels calibrated PMCspectra_calibrated.pdfPMC_MEDMs.png
                      Reply  Tue Apr 18 16:10:00 2017, gautam, Update, PSL, PMC OLTF measured, DAQ channels calibrated PMC_OLTF_170418.pdf
                         Reply  Tue Apr 18 23:37:15 2017, rana, Update, PSL, PMC OLTF measured, DAQ channels calibrated 
                         Reply  Wed Apr 19 15:13:30 2017, gautam, Update, PSL, PMC/MCL multicoherence PMC_MCL_multicoherence.pdf
Message ID: 12944     Entry time: Tue Apr 18 01:01:03 2017     In reply to: 12940     Reply to this: 12945
Author: gautam 
Type: Update 
Category: PSL 
Subject: PMC OLTF measured, DAQ channels calibrated 

Quick entry, details to follow in the AM tomorrow.

  • I calibrated the PMC DAQ channels into physical units - there now exists in the filter modules  cts2m and cts2Hz filter modules, though of course only one must be used at a time
  • Finally measured the PMC OLTF, after moving the PMC PDH error signal demodulation off the servo board - I used the same procedure as Koji when he made the modifications to the PMC servo board, I will put up the algebra here tomorrow. Turns out the previously nominal servo gain of +10dB on the MEDM sliders was a little low, the new nominal gain is +20dB, and has been updated on the MEDM screen.

ToDo:

  • Put up the modified schematic on the 40m DCC tree Done April 18 10pm
  • Check calibration by comparing inferred PMC cavity displacement from error point and control point spectra, using the measured OLTF
  • Finish up looking at multicoherence with MCL and various witness channel combinations

   

Attachment 1: PMCspectra_calibrated.pdf  63 kB  | Hide | Hide all
PMCspectra_calibrated.pdf
ELOG V3.1.3-